OpenStudy (anonymous):

Choose the point which is a solution to the system of inequalities: x – 5y > –7 3x + 4y <–8 (it's < or = to) (–2, 4) (4, 5) (3, –8) (–1, –1) Choose the point which is a solution to the system of inequalities: x – 5y > –7 3x + 4y <–8 (it's < or = to) (–2, 4) (4, 5) (3, –8) (–1, –1) @Mathematics

OpenStudy (valpey):

-2-5*4 = -22 which is not greater than -7 4-5*5 = -21 which is not greater than -7 3-5*(-8) = 43 which is greater than -7 3*3+4*(-8) = -23 which is less than -8 -1-5(-1) = 4 which is greater than -7 3(-1) + 4(-1) = -7 which is not less than -8 so only the third answer is correct.

OpenStudy (anonymous):

3,-8